OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 A. x = 0 and x = 4 B. x = −4 and x = 0 C. x = −20 and x = 4 D. No solution

OpenStudy (anonymous):

@HectorH

OpenStudy (anonymous):

First subtract 10 to get \[\left| x - 2 \right| = 2\]

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

Then what 2 numbers can you plug in so that it makes sense?

OpenStudy (anonymous):

wait is it B?

OpenStudy (anonymous):

No. -4 - 2 = -6

OpenStudy (anonymous):

oh wait lol A

OpenStudy (anonymous):

Correct. Good job :).

OpenStudy (anonymous):

thnx :)